OpenStudy (anonymous):

PLEASE HELP!!! A coin is tossed 8 times. Find the probability P(of exactly 6 occurrences of tails).

OpenStudy (anonymous):

\[\left(\begin{matrix}8 \\ 2\end{matrix}\right)\times( \frac{ 1 }{ 2})^6\times (\frac{ 1 }{ 2 })^2\]

OpenStudy (anonymous):

a. 0.02% c. 0.34% b. 0.17% d. 10.94%

OpenStudy (anonymous):

0.109375 (D)

OpenStudy (anonymous):

Thank you! :)